# Flux4D: Flow-based Unsupervised 4D Reconstruction

### v1

@@ 2 Related Work
Most recently, DrivingRecon lu2024drivingrecon and STORM yang2025storm explore unsupervised generalizable 4D reconstruction for driving scenes, using feed-forward networks to predict the velocities of 3D Gaussians. Despite impressive performance, they can process only sparse (3-4), low-resolution ($\leq 256\times 512$) frames with substantial computational requirements and rely on pre-trained vision models (DeepLabv3+ chen2017rethinking , SAM kirillov2023segment , ViT-Adapter chen2023vision ) for additional supervision, limiting their scalability and applicability. Flux4D achieves better performance with a simpler and more scalable approach, and through our novel incorporation of LiDAR to initialize the scene, can handle full HD images with denser views (>60>60) while being computationally efficient. Please see supp. for more discussions.
@@ 3.4 Improving Realism and Flow
Flux4D-base recovers the overall scene flow accurately (Table 8). We further introduce polynomial motion parameterizations to better model actor behaviors like acceleration, braking or turning. Please see supp. for more details and comparisons. Exploring more advanced velocity models li2025gvfi or implicit flow representations is an exciting direction for future work. To further improve the flow and appearance quality of dynamic actors, we modify the loss function to focus on dynamic regions. Specifically, we render the flow in the image plane and apply pixel-wise re-weighting to the photometric loss. This gives higher importance to faster-moving regions during training, which typically occupy fewer pixels and would contribute less to the overall loss.
@@ 4 Experiments
STORM yang2025storm 17.65 0.367 5.24 0.203 20.79 0.508 4.80 0.238 0.07s
@@ 4.1 Experimental Details
We report standard metrics to measure the photorealism, geometric and motion accuracy using PSNR, SSIM, and depth RMSE (VRMSE) and velocity RMSE (VRMSE). Results are reported on both full images and dynamically moving regions for a comprehensive assessment. For scene-flow quality, we report EPE3D, $Acc_{5}$ and ${Acc}_{10}$ (fraction of points with error $\leq$ 5/10 cm), angular error in radians ($\theta_{\epsilon}$), three-way EPE chodosh2024re : background-static (BS), foreground-static (FS), and foreground-dynamic (FD), bucketed normalized EPE khatri2024can , and inference speed. On WOD, where semantic labels are coarse, we follow EulerFlow vedder2025neural and report bucketed normalized EPE for Background (incl. Signs), Vehicles, Pedestrians, and Cyclists only.
@@ 4.2 Scalable 4D Reconstruction
We further compare Flux4D with SoTA generalizable methods on WOD in Table 4, where we follow the setup in lu2024drivingrecon . The baseline results are from DrivingRecon lu2024drivingrecon paper and we confirmed the setup and results with the authors to ensure accurate comparison. Flux4D surpasses DrivingRecon by +5.99 dB in PSNR and +0.21 in SSIM, demonstrating its effectiveness for unsupervised dynamic scene reconstruction. Please see supp. for qualitative comparisons.
@@
While Flux4D primarily focuses on reconstruction and is not specifically designed for scene flow estimation, we further evaluate its performance on PandaSet compared with representative scene flow estimation methods using standard scene flow metrics in Table 5 and 6. Please see supp. for comparisons on WOD. Although not designed for scene flow estimation, Flux4D achieves superior performance across most scene flow metrics using only reconstruction-based supervision (RGB + depth). Notably, it outperforms other methods on smaller or less common object categories such as wheeled VRUs, other vehicles, and pedestrians, as shown in bucketed evaluations. These results highlight a promising path to unifying state-of-the-art scene flow estimation khatri2024can ; khatri2024can ; kim2025flow4d ; li2025uniflow and reconstruction within a single framework.
@@
We evaluate Flux4D’s capability for future frame prediction beyond the observed frames. This challenging task requires precise motion estimation, temporal consistency, occlusion reasoning, and a comprehensive 4D scene understanding. As shown in Table 4, Flux4D outperforms existing unsupervised methods in both photometric accuracy and geometric consistency. Moreover, Flux4D even outperforms supervised approaches that rely on imperfect explicit annotations for extrapolation, demonstrating the robustness of our predicted scene representation and the effectiveness of unsupervised scene flow prediction. This highlights Flux4D’s ability to model scene dynamics, which is critical for world modeling, simulation, and scene understanding in autonomous systems. We report dynamic-only metrics in Table 4 and defer full-image metrics to supp.
@@
Table 8 evaluates Flux4D’s key design components. Iterative refinement significantly enhances image quality and geometric accuracy metrics. Polynomial motion modeling improves motion prediction performance. Table 8 demonstrates that our static-preference prior is essential to learning accurate flow, and that velocity reweighting improves performance on the dynamic elements. Please refer to supp. for qualitative comparisons.
@@
[…]
@@ 5 Limitations
Although Flux4D achieves SoTA 4D reconstruction without any annotations or pre-trained models, three key limitations remain: (1) flow estimation for highly dynamic actors with complex motion patterns is challenging, which could be mitigated by leveraging larger and more diverse training data; (2) iterative approach for long-horizon reconstruction creates visible inconsistencies at transition points; and (3) the method assumes a simple pinhole camera model with clean LiDAR data, limiting applicability with rolling shutter cameras or noisy sensor inputs. Please see supp. for more examples. Future work will focus on scaling to larger datasets, developing a unified temporal representation for seamless long-term reconstruction, and improving robustness to real-world sensor imperfections. Furthermore, Flux4D’s explicit 3D representation offers interpretable structure for world models. Overall, we believe that our simple and scalable design serves as a foundation for the community to build upon, enabling further advancements in 4D reconstruction.

### v2

@@ 2 Related Work
Most recently, DrivingRecon lu2024drivingrecon and STORM yang2025storm explore unsupervised generalizable 4D reconstruction for driving scenes, using feed-forward networks to predict the velocities of 3D Gaussians. Despite impressive performance, they can process only sparse (3-4), low-resolution ($\leq 256\times 512$) frames with substantial computational requirements and rely on pre-trained vision models (DeepLabv3+ chen2017rethinking , SAM kirillov2023segment , ViT-Adapter chen2023vision ) for additional supervision, limiting their scalability and applicability. Flux4D achieves better performance with a simpler and more scalable approach, and through our novel incorporation of LiDAR to initialize the scene, can handle full HD images with denser views (>60>60) while being computationally efficient. Please see Appendix E for more discussions.
@@ 3.4 Improving Realism and Flow
Flux4D-base recovers the overall scene flow accurately (Table 8). We further introduce polynomial motion parameterizations to better model actor behaviors like acceleration, braking or turning. Please see Appendix A.1 for more details and comparisons. Exploring more advanced velocity models li2025gvfi or implicit flow representations is an exciting direction for future work. To further improve the flow and appearance quality of dynamic actors, we modify the loss function to focus on dynamic regions. Specifically, we render the flow in the image plane and apply pixel-wise re-weighting to the photometric loss. This gives higher importance to faster-moving regions during training, which typically occupy fewer pixels and would contribute less to the overall loss.
@@ 4 Experiments
STORM yang2025storm 17.65 0.367 5.24 0.203 20.79 0.508 4.80 0.238 0.08s
Flux4D-base (Ours) 18.62 0.454 2.06 0.181 20.88 0.575 1.37 0.211 0.13s
@@ 4.1 Experimental Details
We report standard metrics to measure the photorealism, geometric and motion accuracy using PSNR, SSIM, and depth MAE (DMAE) and normalized velocity RMSE (VRMSE). Please see Appendix C for more details. Results are reported on both full images (non-sky regions) and dynamically moving regions. For scene-flow quality, we report EPE3D, $Acc_{5}$ and ${Acc}_{10}$ (fraction of points with error $\leq$ 5/10 cm), angular error in radians ($\theta_{\epsilon}$), three-way EPE chodosh2024re : background-static (BS), foreground-static (FS), and foreground-dynamic (FD), bucketed normalized EPE khatri2024can , and inference speed. On WOD, where semantic labels are coarse, we follow EulerFlow vedder2025neural and report bucketed normalized EPE for Background (incl. Signs), Vehicles, Pedestrians, and Cyclists only.
@@ 4.2 Scalable 4D Reconstruction
We further compare Flux4D with SoTA generalizable methods on WOD in Table 4, where we follow the setup in lu2024drivingrecon . The baseline results are from DrivingRecon lu2024drivingrecon paper and we confirmed the setup and results with the authors to ensure accurate comparison. Flux4D surpasses DrivingRecon by +5.99 dB in PSNR and +0.21 in SSIM, demonstrating its effectiveness for unsupervised dynamic scene reconstruction. Please see Appendix D for qualitative comparisons.
@@
While Flux4D primarily focuses on reconstruction and is not specifically designed for scene flow estimation, we further evaluate its performance on PandaSet compared with representative scene flow estimation methods using standard scene flow metrics in Table 5 and 6. Please see Appendix D for comparisons on WOD. Although not designed for scene flow estimation, Flux4D achieves superior performance across most scene flow metrics using only reconstruction-based supervision (RGB + depth). Notably, it outperforms other methods on smaller or less common object categories such as wheeled VRUs, other vehicles, and pedestrians, as shown in bucketed evaluations. These results highlight a promising path to unifying state-of-the-art scene flow estimation khatri2024can ; kim2025flow4d ; li2025uniflow and reconstruction within a single framework.
@@
We evaluate Flux4D’s capability for future frame prediction beyond the observed frames. This challenging task requires precise motion estimation, temporal consistency, occlusion reasoning, and a comprehensive 4D scene understanding. As shown in Table 4, Flux4D outperforms existing unsupervised methods in both photometric accuracy and geometric consistency. Moreover, Flux4D even outperforms supervised approaches that rely on imperfect explicit annotations for extrapolation, demonstrating the robustness of our predicted scene representation and the effectiveness of unsupervised scene flow prediction. This highlights Flux4D’s ability to model scene dynamics, which is critical for world modeling, simulation, and scene understanding in autonomous systems. We report dynamic-only metrics in Table 4 and defer full-image metrics to Appendix D.
@@
Table 8 evaluates Flux4D’s key design components. Iterative refinement and polynomial motion significantly enhance image quality, geometric accuracy and motion prediction performance. Table 8 demonstrates that our static-preference prior is essential to learning accurate flow, and that velocity reweighting improves performance on the dynamic elements. Please refer to Appendix D for qualitative comparisons.
@@
PSNR $\uparrow$ SSIM $\uparrow$ $D_{\mathrm{MAE}}\downarrow$ $V_{\mathrm{RMSE}}\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ $D_{\mathrm{MAE}}\downarrow$ $V_{\mathrm{RMSE}}\downarrow$ EPE-3way $\downarrow$
@@ 5 Limitations
Although Flux4D achieves SoTA 4D reconstruction without any annotations or pre-trained models, three key limitations remain: (1) flow estimation for highly dynamic actors with complex motion patterns is challenging, which could be mitigated by leveraging larger and more diverse training data; (2) iterative approach for long-horizon reconstruction creates visible inconsistencies at transition points; and (3) the method assumes a simple pinhole camera model with clean LiDAR data, limiting applicability with rolling shutter cameras or noisy sensor inputs. Please see Appendix F for more examples. Future work will focus on scaling to larger datasets, developing a unified temporal representation for seamless long-term reconstruction, and improving robustness to real-world sensor imperfections. Furthermore, Flux4D’s explicit 3D representation offers interpretable structure for world models. Overall, we believe that our simple and scalable design serves as a foundation for the community to build upon, enabling further advancements in 4D reconstruction.